OpenStudy (anonymous):

5z^3-z^4>or equal 0 solve for z

OpenStudy (anonymous):

\[5z^3-z^4\ge0\]

hartnn (hartnn):

can u take z^3 common from those 2 terms and tell me what u get?

OpenStudy (anonymous):

z^3(5-z)

hartnn (hartnn):

z^3(5-z)>=0 so that i can write as z^3>=0 , 5-z>=0 got this?

OpenStudy (anonymous):

so 0<z<5?

hartnn (hartnn):

absolutely,just put = sign also with <

OpenStudy (anonymous):

\[0 \le z \le 5\]

hartnn (hartnn):

yup :)